An automobile accelerates from a speed of 32.2 km/h to 96.6 km/h in a distance of 89.3

m, which requires 5 s.

(a) If the acceleration is constant, find the acceleration in meters per second squared.

(b) Same as part (a) except the acceleration is not constant. What is the average

acceleration in meters per second squared?


1
Expert's answer
2021-11-06T01:35:29-0400

(a)


"a=\frac{v_2^2-v_1^2}{2d}=3.58\text{ m/s}^2."

(b)


"a=\frac{v_2-v_1}{t}=3.58\text{ m/s}^2."